OpenStudy (eturpin1):

Before a nuclear reaction, there is 0.01000 kg of uranium. After the nuclear reaction, there is 0.00950 kg of uranium. How much energy was released? (h = 6.63 × 1034 Js)

OpenStudy (michele_laino):

here we have to apply the Einstein's relationship between mass and energy, namely: \[\Large \Delta E = \Delta m \cdot {c^2}\] where: \[\Large \Delta m = 0.01 - 0.0095 = ...{\text{Kg}}\]

OpenStudy (eturpin1):

5x10^-4?

OpenStudy (michele_laino):

\[\Large \Delta E = \Delta m \cdot {c^2} = 0.0045 \times 9 \times {10^{16}} = ...{\text{Joules}}\]

OpenStudy (eturpin1):

so 4.50x10^13 J?

OpenStudy (michele_laino):

that's right!

OpenStudy (eturpin1):

thanks so much! Could you help me with a couple more if you dont mind?

OpenStudy (michele_laino):

sorry I have made a typo the right formula is: \[\Large \Delta E = \Delta m \cdot {c^2} = 0.0005 \times 9 \times {10^{16}} = 4.5 \times {10^{13}}\;{\text{Joules}}\]

OpenStudy (michele_laino):

ok! I will help you!

OpenStudy (eturpin1):

Thanks so much! Which of the following processes is used by NASA to make electric power on deep space probes? radioactive decay fission fusion X-ray bombardment

OpenStudy (michele_laino):

I'm sorry I don't know that answer, since I'm Italian, and I never studied, at my university the technologies used by NASA

OpenStudy (eturpin1):

oh okay, could you help me with another one with some calculating values?

OpenStudy (michele_laino):

ok!

OpenStudy (eturpin1):

What is the energy of a single photon of red light with a wavelength of 700.0 nm? Remember that E = hf and Planck's constant is expressed by h = 6.63 × 1034 Js = 4.14 × 1015 eVs. 2.84 × 1019 J 4.98 × 1019 J 2.84 × 1014 J 4.29 × 1014 J

OpenStudy (michele_laino):

we can write: \[\lambda \nu = c\] where \lambda is the wavelenght of your radiation \nu is its frequency, and c is the light speed so we have: \[\nu = \frac{c}{\lambda }\]

OpenStudy (michele_laino):

now the anergy of a photon, which is the mediator particle of the electromagnetic field, is given by the subsequent expression: \[\Large E = h\nu = \frac{{hc}}{\lambda }\]

OpenStudy (michele_laino):

so you have to substitute your data into the above formula, what do you get?

OpenStudy (eturpin1):

Are we looking for c?

OpenStudy (michele_laino):

I got this: \[\Large E = h\nu = \frac{{hc}}{\lambda } = 2.84 \times {10^{ - 19}}Joules\]

OpenStudy (michele_laino):

keep in mind that: \[\Large h = 6.62 \times {10^{ - 34}}Joules \times \sec \]

OpenStudy (eturpin1):

so do we multiply the two?

OpenStudy (michele_laino):

here are more steps: \[\Large E = h\nu = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{700 \times {{10}^{ - 9}}}}\]

OpenStudy (eturpin1):

so 2.85x10^-37?

OpenStudy (michele_laino):

no, 2.84*10^(-19)

OpenStudy (eturpin1):

hmm, i dont know how I got that...

OpenStudy (michele_laino):

hint: \[700nm = 700 \times {10^{ - 9}}meters = 7 \times {10^{ - 7}}meters\]

OpenStudy (eturpin1):

ohh okay, I put it in incorrectly! do you have time for another?

OpenStudy (michele_laino):

yes!

OpenStudy (eturpin1):

A paleontologist measures the ratio of carbon-14 to carbon-12 in a fossil skull found at a site. What technique is he using and for what purpose? (Points : 3) using alpha emission to determine the structure of the skull using radioactive decay to date the skull using radioactive decay to determine the structure of the skull using gamma emissions to date the skull

OpenStudy (michele_laino):

The ratio C14 to C12 is used to get the date of a sample which contains carbon, and the radioactive law is employed to get that date from that sample

OpenStudy (michele_laino):

more precisely, using the radioactive law, we can get when a sample has ceased to live

OpenStudy (michele_laino):

so, what is the right option?

OpenStudy (eturpin1):

using radioactive decay to date the skull?